OpenStudy (anonymous):

find the derivative of y=x(e^2)-e

jimthompson5910 (jim_thompson5910):

It's \[\Large y = x(e^2)-e\] right?

OpenStudy (anonymous):

yes!

jimthompson5910 (jim_thompson5910):

e is just a number (it's roughly 2.71828...) it's like pi

jimthompson5910 (jim_thompson5910):

so e^2 is just a number

jimthompson5910 (jim_thompson5910):

so e^2 is effectively the coefficient of the x term

jimthompson5910 (jim_thompson5910):

so if I asked you to derive y = 2x+1 for me, what would you say?

OpenStudy (anonymous):

2

jimthompson5910 (jim_thompson5910):

good, if you have a line, the derivative is just the slope of that line

OpenStudy (anonymous):

I know the answer to the problem, I just don't see how to get to it

jimthompson5910 (jim_thompson5910):

this is no different (it's just a bit more convoluted)

jimthompson5910 (jim_thompson5910):

the slope in this case e^2, so the derivative of y is e^2

OpenStudy (anonymous):

they say it's dy/dx = e^2 - e^x

jimthompson5910 (jim_thompson5910):

oh then there has to be something more to the problem

jimthompson5910 (jim_thompson5910):

so the problem can't be \[\Large y = x(e^2)-e\]

jimthompson5910 (jim_thompson5910):

If the problem was \[\Large y = x(e^2)-e^x\] then that would the correct answer

OpenStudy (anonymous):

oh, it's that, yes, but how does that work :(

OpenStudy (anonymous):

i'm so confused

jimthompson5910 (jim_thompson5910):

you see how the derivative of x(e^2) is e^2 right?

OpenStudy (anonymous):

yes!

jimthompson5910 (jim_thompson5910):

so the only thing left is to take the derivative of e^x this will give you e^x

jimthompson5910 (jim_thompson5910):

this is why e is so special

OpenStudy (anonymous):

OH, ok thank you! can I chat you if I have more questions?

jimthompson5910 (jim_thompson5910):

it's because it is its own derivative

jimthompson5910 (jim_thompson5910):

sure

jimthompson5910 (jim_thompson5910):

and I'm sure you have learned this rule if y = b^x then dy/dx = b^x*ln(b) if b = e, then you get y = e^x dy/dx = e^x

OpenStudy (anonymous):

because my whole work for the night is going to be derivatives of logs, ln, and e

OpenStudy (anonymous):

we have all the rules, lnx = 1/x, lnu = 1/u du/dx, e^x = e^x, logau = 1/xlna, etc

jimthompson5910 (jim_thompson5910):

yes, if y = ln(x) then dy/dx = 1/x etc etc

OpenStudy (anonymous):

I have a question

OpenStudy (anonymous):

there's a difference between logau and logax

OpenStudy (anonymous):

a being the base

OpenStudy (anonymous):

when do you use each formula, you use logax when there's an x as the variable right?

jimthompson5910 (jim_thompson5910):

what were the formulas you are referring to again?

jimthompson5910 (jim_thompson5910):

if y = log(a,x) then dy/dx = 1/(x*ln(a)) right?

jimthompson5910 (jim_thompson5910):

log(a,x) means "log base 'a' of x"

OpenStudy (anonymous):

i mean log base 'a' of x

OpenStudy (anonymous):

but when do you use log base a of x and log base a of u?

jimthompson5910 (jim_thompson5910):

i know

jimthompson5910 (jim_thompson5910):

unless i can see what formulas you're working with, they both sound the same (just replace x with u)...so I see no difference

OpenStudy (anonymous):



jimthompson5910 (jim_thompson5910):

i think you mean log(a,u) = 1/u*ln(a) * dy/dx

jimthompson5910 (jim_thompson5910):

or to be more accurate the derivative of log(a,u) with respect to x is 1/( u*ln(a) ) * du/dx

OpenStudy (anonymous):

then what did he mean with log (x, u) equaling that without du/dx?

jimthompson5910 (jim_thompson5910):

log(x, u) means log base x of u shouldn't it be log(a, u) or log(a, x)

OpenStudy (anonymous):

i meant, log (a, x)

OpenStudy (anonymous):

seriously having a mental crash right now, it's all jumbling, nothings making sense, and the tests tomorrow

jimthompson5910 (jim_thompson5910):

have you heard of the chain rule?

OpenStudy (anonymous):

yeah!

jimthompson5910 (jim_thompson5910):

that's what's going on normally, the derivative of log(a,x) is just \[\Large \frac{1}{x\ln(a)}\]

jimthompson5910 (jim_thompson5910):

but when you introduce the 'u' in there, it complicates things because you have to use the chain rule

jimthompson5910 (jim_thompson5910):

does that make sense?

OpenStudy (anonymous):

what's the difference between a, and u?

OpenStudy (anonymous):

*x and u

OpenStudy (anonymous):

when it says x does it mean, literally just the variable x

OpenStudy (anonymous):

or could you put log 4 x^2 and stilll use log a x

OpenStudy (anonymous):

or would the x^2 be used as u

jimthompson5910 (jim_thompson5910):

u and x are their own variables, but when we say dy/dx we mean "derive the function y with respect to x" So... if you have something with u in it (and no x), then either u is assumed to be a constant (not likely though) or it's a function of x on its own ex: if u = 2x and y = log(3, u) this is the same thing as saying y = log(3, 2x)

jimthompson5910 (jim_thompson5910):

this is why it's a bit more complicated

OpenStudy (anonymous):

so log 5 x you would use log a x

OpenStudy (anonymous):

or something with a constant base, and x

OpenStudy (anonymous):

but you use log a u for anything pretty much?

OpenStudy (anonymous):

brb going to shower, to clear the headache

jimthompson5910 (jim_thompson5910):

yes, log(a, x) allows you to say log(5, x), but you can't stray from that (so only 'a' is allowed to change) but log(a, u) can be more difference since you can replace 'u' with any expression of x

jimthompson5910 (jim_thompson5910):

different*